OpenStudy (anonymous):

Fiona has $18 to spend. She spent $4.25, including tax, to buy a notebook. She needs to save $9.75, but she wants to buy a snack. If crackers cost $0.50 per package including tax, what inequality would show the maximum number of packages that Fiona can buy? Solve your inequality by showing your work and steps.

OpenStudy (anonymous):

@KennyVanGemert

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

@heavenfun13 do you think you can help

OpenStudy (anonymous):

hmmpft im not good at math im so sorry id love to help but i cant

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

let me try

OpenStudy (danjs):

starts with 18 spends 4.25 18 - 4.25 spends .50 on each cracker 18 - 4.25 - 0.50*(#crackers) all that should be greater than or equal to 9.75, the amount she needs to keep

OpenStudy (anonymous):

I need to make an inequality and then solve it

OpenStudy (anonymous):

AM I ALONE AGAIN I HAVE BEEN WAITING FOR OVER AN HOUR

OpenStudy (anonymous):

@camerondoherty

OpenStudy (anonymous):

@camerondoherty may you please help me.

OpenStudy (danjs):

i just told you the inequlity

OpenStudy (danjs):

18 - 4.25 - 0.50*(#crackers) is greater than or equal to 9.75 solve for (#crackers)

OpenStudy (anonymous):

doesnt there need to be a variable

OpenStudy (danjs):

starts with 18 spends 4.25 left with 13.75 spends 0.50 per crackers ---0.50*(#crackers) wants at least 9.75 left over 13.75 - 0.50*(#crackers) >= 9.75

OpenStudy (danjs):

\[13.75 - 9.75 \ge 0.5*(crackers)\]

OpenStudy (danjs):

\[4 \ge 0.5*(crackers)\] \[8 \ge crackers\] the amount of crackers she can get has to be less than or equal to 8

OpenStudy (anonymous):

I mostly need the steps to solve it

OpenStudy (danjs):

13.75 - 0.50*(#crackers) >= 9.75 add 0.50*(crackers) to each side 13.75 >= 9.75 + 0.5*(crackers) subtract 9.75 from both sides 4 >= 0.5*(crackers) divide by 0.5 on both sides 8 >= (crackers) eight is greater than or equal to the amount of crackers